OpenStudy (anonymous):

What is the estimate for 8 1/8 * 40?

OpenStudy (anonymous):

8*(1/8) = what

OpenStudy (anonymous):

65

OpenStudy (anonymous):

no its just 1

OpenStudy (anonymous):

stop messing around

hartnn (hartnn):

8 1/8 is mixed fraction ? @MauriceLong

OpenStudy (anonymous):

the 8's cancel so its 1*40 =

OpenStudy (anonymous):

oh ok i get it now.. its a mixed fraction

OpenStudy (anonymous):

it becomes 40

hartnn (hartnn):

@ilikephysics2 please don't be rude

hartnn (hartnn):

yup, so 8 1/8 = 65/8 and when multiplied by 40 it gives 65/8 * 40 could u solve this ?

OpenStudy (anonymous):

do you put a 1 under 40 to make it a fraction

hartnn (hartnn):

nopes, \(\huge\frac{65}{8}.40= 65*5=?\)

hartnn (hartnn):

because 8*5=40

OpenStudy (anonymous):

65*5 = 325

OpenStudy (anonymous):

if its a mixed fraction then its 65/8 * 40 =

OpenStudy (anonymous):

325

hartnn (hartnn):

yup, 325 is correct!